OpenStudy (anonymous):

solve for x. answer should not be approximation: 1+3ln(x-2)=13 and 2(5)^4x=7(5)^x is there a way to put it into steps to make it easier to see how to set these up

OpenStudy (anonymous):

1+3ln(x-2)=13 3ln(x-2)=12 ln(x-2)=4 e^(ln(x-2))=e^4 x-2=e^4 x=e^4-2

OpenStudy (anonymous):

Do you have anything for the other part of the question? I know it's probably easy to solve but for some reason I can't seem to make it work for me, if you have anything I'd really appreciate it. Thanks

OpenStudy (turingtest):

\[2(5)^{4x}=7(5)^x\implies5^{3x}=\frac72\]can you see what to do now?

OpenStudy (anonymous):

I think I do 2(5)\[2(5)^{4x}\div(5)^{x} =7 \]\[2(5)^{x}\div2=7\div2\]\[5^{x}=\frac{ 7 }{ 2 }\] Do these steps look right or am I just going about it the wrong way?